5.6 Assessment Task




Directions. Answer the following items. You may use online calculators and solvers in answering.




1. For each of the following data sets determine the mean:




a. 72, 14,8, 11,57,54,31, 11,67,11,19,3,66 b. 63,9,87,16, 2, 96,13,67,34




c. 1,6,8,2,7,2,9,4,8,9,8,6 d. 12, 0,4,5,8, 3,6, 35,47




e. 2, 7, 17, 33, 67,73, 88, 33, 92, 57,33




2. For each of the following data sets determine the median:




a. 48, 78, 10, 66, 45, 57,96,67, 40, 66, 63,8, 20 b. 28, 3, 10, 60, 8, 23, 45, 97,11, 10




c. 6,9,1, 6, 4, 8, 1, 7, 8, 3, 1, 0




d. 63, 9, 86, 16, 2, 97,24, 67, 34,40 e. 8, 4, 64, 99, 11, 42, 15,88, 54, 77,42




3. For each of the following data sets determine the mode:




a. 98, 37,5,33, 96, 67, 43, 33, 91, 33,32,8,11 b. 104, 2, 51,31,8, 101, 104, 18,47




c. 3, 4, 5, 3,9,8,5,7,2,5,1




d. 19, 1,9,6,4,2,13,15,24,2




e. 8,9,39,44,55,90,19,44,28,69,44




43/76

1
Expert's answer
2022-05-06T12:39:50-0400

1.

a.


"mean=\dfrac{1}{13}(98+37+5+33+96+67"

"+43+33+91+33+32+8+11)=\dfrac{587}{13}"

b.


"mean=\dfrac{1}{9}(63+9+87+16+2+96+13"

"+67+34)=43"

c.


"mean=\dfrac{1}{12}(1+6+8+2+7+2+9"

"+4+8+9+8+6)=\dfrac{35}{6}"

d.


"mean=\dfrac{1}{9}(12+0+4+5+8+3+6+35+47)"

"=\dfrac{40}{3}"

e.


"mean=\dfrac{1}{11}(2+7+17+33+67+73"

"+88+33+92+57+33)=\dfrac{502}{11}"

2.

a.


"8,10,20,40,45,48, 57,63,66,66,67,78, 96"

"median=57"

b.


"3,8,10,10,11, 23, 28, 45, 60, 97"

"median=\dfrac{11+23}{2}=17"

c.


"0,1,1,1,3,4,6,6,7,8,8,9"




"median=\dfrac{4+6}{2}=5"

d.


"2,9,16, 24,34,40, 63, 67, 86, 97"

"median=\dfrac{34+40}{2}=37"


e.


"4,8, 11,15, 42, 42, 54, 64, 77, 88, 99"

"median=42"

3.

a.


"5,8,11,32, 33,33, 33, 37, 43, 67, 91, 96, 98"

"mode=33\ appeared\ 3\ times"

b.


"2,8, 18, 31,47, 51, 101, 104, 104"

"mode=104\ appeared\ 2\ times"

c.


"1, 2, 3, 3, 4, 5, 5, 5, 7, 8, 9"

"mode=5\ appeared\ 3\ times"

d.


"1,2,2,4, 6, 9, 13, 15,19, 24"

"mode=2\ appeared\ 2\ times"

e.


"8,9, 19, 28, 39,44, 44, 44, 55, 69, 90"




"mode=44\ appeared\ 3\ times"